# Hybrid estimation for a mixed fractional Black-Scholes model with random effects from discrete time observations

### v3

@@ 7.7 Proof of Proposition 5
For each $i=1,\ldots,N$, the random effect is expressed by $\phi_{i}=\theta_{i}+\frac{1}{2}\sigma^{2}.$ Then, by a simple plug-in, an estimator of $\phi_{i}$ is given by $\widehat{\phi}_{i}:=\widehat{\theta}_{i}+\frac{1}{2}\widehat{\sigma}^{2},$ where $\theta_{i}$ is given by (3.2) and $\hat{\sigma}^{2}$ is as expressed in (3.10). From Proposition 1, we already know that $\widehat{\theta}_{i}\xrightarrow{a.s}\theta_{i}$, as $n\to\infty$ and from Proposition 4, $\widehat{\sigma}^{2}\xrightarrow{a.s.}\sigma^{2},$ as $N\to\infty$. Combining the two results, we conclude that
$\widehat{\theta}_{i}+\frac{1}{2}\widehat{\sigma}^{2}\xrightarrow{a.s.}\theta_{i}+\frac{1}{2}\sigma^{2},\quad\text{as}\quad n,N\to\infty.$
@@
Now we proceed to prove the asymptotic normality of $\hat{\phi}_{i}$ stated in the second assertion of Proposition 5. We first decompose the estimation error for each $i$, as follows
@@
As we can observe, the estimation error (7.11) involves both the subject-specific term $\widehat{\theta}_{i}-\theta_{i}$, which depends on the number of observations $n$ and the global parameter term $\widehat{\sigma}^{2}-\sigma^{2}$, which depends on the number of subjects $N$. Therefore, the asymptotic distribution of (7.11) depends on the relative growth of $n$ and $N$. For this reason, we distinguish two asymptotic regimes: in the first, we let $n\to\infty$ first and then $N\to\infty$, and in the second, we let $n,N\to\infty$ jointly.
Case 1: $n\to\infty$, then $N\to\infty$
We first fix $N$ and assume that $n\to\infty$. For each subject $i=1,\ldots,N$, $\hat{\theta}_{i}$ is strongly consistent. Consequently, $\widehat{\theta}_{i}-\theta_{i}\xrightarrow{\mathbb{P}}0,\!\ \mathrm{as}\!\ \ n\to\infty$. We now let $N\to\infty$. From Proposition 3, we have
$\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})\xrightarrow{d}\mathcal{N}(0,v_{\sigma}^{2}),$
where $v_{\sigma}^{2}$ is the asymptotic variance of $\widehat{\sigma}^{2}$. Combining the two convergences via Slutsky’s theorem, we deduce that as $n\to\infty$ first and then $N\to\infty$, $\hat{\phi}_{i}$ is asymptotically normal and satisfies
$\sqrt{N}(\widehat{\phi}_{i}-\phi_{i})\xrightarrow{d}\mathcal{N}\!\Big(0,\tfrac{1}{4}v_{\sigma}^{2}\Big).$
Case 2: $n,N\to\infty$ simultaneously
It is already known from Proposition 1 that the scaling term for $\widehat{\theta}_{i}-\theta_{i}$ is $n^{1-H}$ while from Proposition 3, the scaling term for $\widehat{\sigma}^{2}-\sigma^{2}$ is $\sqrt{N}$. Scaling the estimation error (7.11) by $n^{1-H}$, we obtain
$n^{1-H}(\widehat{\phi}_{i}-\phi_{i})=n^{1-H}\Big(\widehat{\theta}_{i}-\theta_{i}\Big)+\dfrac{1}{2}\frac{n^{1-H}}{\sqrt{N}}\sqrt{N}\Big(\widehat{\sigma}^{2}-\sigma^{2}\Big).$ (7.12)
For the first term of the right-hand side of (7.12), Proposition 1 ensures that
$n^{1-H}(\widehat{\theta}_{i}-\theta_{i})\xrightarrow{d}\mathcal{N}\Big(0,v_{\theta}^{2}\Big)$, where $v_{\theta}^{2}:=\gamma^{2}h^{2H-2}$ is the asymptotic variance of $n^{1-H}(\widehat{\theta}_{i}-\theta_{i})$. For the second term of the right-hand side of (7.12), we should control the ratio $\dfrac{n^{1-H}}{\sqrt{N}}$.
We distinguish three sub-cases:
  • If $\dfrac{n^{1-H}}{\sqrt{N}}\to c\in(0,\infty)$, then both terms in (7.11) contribute. Furthermore, since $\widehat{\theta}_{i}$ is constructed from subject $i$ and $\widehat{\sigma}^{2}$ is calculated from subjects averaged by $N$, the contribution of $\theta_{i}$ becomes negligible as $N\to\infty$, which ensures the asymptotic independence of $\widehat{\theta}_{i}-\theta_{i}$ and $\widehat{\sigma}^{2}-\sigma^{2}$. Hence, the limit of the error term (7.11) is Gaussian with variance $v_{\theta_{i}}^{2}+\frac{c^{2}}{4}v_{\sigma}^{2}$, where $v_{\sigma}^{2}$ is the asymptotic variance of $\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})$, and we obtain
    $n^{1-H}(\widehat{\phi}_{i}-\phi_{i})\xrightarrow{d}\mathcal{N}\!\Big(0,v_{\theta}^{2}+\frac{c^{2}}{4}v_{\sigma}^{2}\Big).$
  • If $\dfrac{n^{1-H}}{\sqrt{N}}\to 0$, the term $\widehat{\sigma}^{2}-\sigma^{2}$ vanishes and we deduce that
    $n^{1-H}(\widehat{\phi}_{i}-\phi_{i})\xrightarrow{d}\mathcal{N}\!\Big(0,v_{\theta}^{2}\Big).$
  • If $\dfrac{n^{1-H}}{\sqrt{N}}\to\infty$, the term $\widehat{\theta}_{i}-\theta_{i}$ vanishes and the estimation error is dominated by the term $(\widehat{\sigma}^{2}-\sigma^{2})$ and the suitable normalization is $\sqrt{N}$. Hence,
    $\sqrt{N}(\widehat{\phi}_{i}-\phi_{i})\xrightarrow{d}\mathcal{N}\!\Big(0,v_{\sigma}^{2}\Big).$
This completes the proof of Proposition 5.
The remainder of this section is devoted to the proofs of the asymptotic properties of the distribution function estimator $\hat{F}_{m,n,N}$. To that aim, we introduce the following lemma which will be needed later.
@@ 7.8 Proof of Proposition 6
For all $x\in[-1,1]$, the expectation of the estimator $\hat{F}_{m,n,N}$ is given by
$\mathbb{E}\left(\hat{F}_{m,n,N}(x)\right)$ $=\sum\limits_{j=1}^{m}\frac{1}{N}\sum\limits_{i=1}^{N}\mathbb{E}\left(\mathds{1}_{\left\{\widehat{\phi}_{i}\leq x_{j}\right\}}\right)\mathcal{L}_{j}(x)$
$=\sum\limits_{j=1}^{m}\mathbb{P}\left(\hat{\phi}_{1}\leq x_{j}\right)\mathcal{L}_{j}(x).$
@@
$\underset{n,N\to\infty}{\lim}\mathbb{P}\left(\hat{\phi}_{1}\leq x_{j}\right)=\mathbb{P}\left(\phi_{1}\leq x_{j}\right).$ (7.13)
@@
$\lim\limits_{n,N\to\infty}\mathbb{E}\left(\hat{F}_{m,n,N}(x)\right)=\sum\limits_{j=1}^{m}\mathbb{P}\left(\phi_{1}\leq x_{j}\right)\mathcal{L}_{j}(x)=\sum\limits_{j=1}^{m}F(x_{j})\mathcal{L}_{j}(x).$ (7.14)
@@
$\lim\limits_{n,N\to\infty}\mathbb{E}\left(\hat{F}_{m,n,N}(x)\right)$ $=F(x)+f(x)\sum\limits_{j=1}^{m}\left(x_{j}-x\right)\mathcal{L}_{j}(x)+\frac{1}{2}f^{\prime}(x)\sum\limits_{j=1}^{m}\left(x_{j}-x\right)^{2}\mathcal{L}_{j}(x)$
@@
$\lim\limits_{n,N\to\infty}\mathbb{E}\left(\hat{F}_{m,n,N}(x)\right)=F(x)+\frac{\pi}{m^{2}}T_{m}(x)\left[\frac{1}{4}\left(x-1\right)f^{\prime}(x)-\frac{1}{2}f(x)\right]+O(m^{-2}),$
@@
Let us now focus on calculating the variance of our estimator presented by the second assertion of Proposition 6. For all $x\in[-1,1]$, we have
$\hat{F}_{m,n,N}(x)-\mathbb{E}\left(\hat{F}_{m,n,N}(x)\right)$ $=\sum\limits_{j=1}^{m}\hat{F}_{N}\left(x_{j}\right)\mathcal{L}_{j}(x)-\sum\limits_{j=1}^{m}\mathbb{P}\left(\widehat{\phi}_{i}\leq x_{j}\right)\mathcal{L}_{j}(x)$
$=\frac{1}{N}\sum\limits_{j=1}^{m}\left[\sum\limits_{i=1}^{n}\left(\mathds{1}_{\left\{\widehat{\phi}_{i}\leq x_{j}\right\}}-\mathbb{P}\left(\widehat{\phi}_{i}\leq x_{j}\right)\right)\right]\mathcal{L}_{j}(x)$
$=\frac{1}{N}\sum\limits_{i=1}^{N}Z_{i,N}$
where $Z_{i,N}:=\sum\limits_{j=1}^{m}\left(\mathds{1}_{\left\{\widehat{\phi}_{i}\leq x_{j}\right\}}-\mathbb{P}\left(\widehat{\phi}_{i}\leq x_{j}\right)\right)\mathcal{L}_{j}(x).$ Note that $Z_{i,N}$ are i.i.d. with mean zero. It follows that
$Var\left(\hat{F}_{m,n,N}(x)\right)$ $=\mathbb{E}\left[\left(\hat{F}_{m,n,N}(x)-\mathbb{E}\left(\hat{F}_{m,n,N}(x)\right)\right)^{2}\right]$
$=\frac{1}{N^{2}}\mathbb{E}\left(\sum\limits_{i=1}^{N}Z_{i,N}^{2}\right)$
$=\frac{1}{N}\mathbb{E}\left(Z_{1,N}^{2}\right).$
For $i=1,\ldots,N$ and $j=1,\ldots,m$, we define $H_{1}(x_{j}):=\mathds{1}_{\left\{\widehat{\phi}_{i}\leq x_{j}\right\}}-\mathbb{P}\left(\widehat{\phi}_{i}\leq x_{j}\right).$ Then,
$\mathbb{E}\left(Z_{1,N}^{2}\right)$ $=\mathbb{E}\left[\left(\sum\limits_{j=1}^{m}H_{1}(x_{j})\mathcal{L}_{j}\left(x\right)\right)^{2}\right]$
$=\mathbb{E}\left[\sum\limits_{j,l=1}^{m}H_{1}(x_{j})\mathcal{L}_{j}\left(x\right)H_{1}(x_{l})\mathcal{L}_{l}\left(x\right)\right]$
$=\sum\limits_{j,l=1}^{m}\mathbb{E}\left[H_{1}(x_{j})H_{1}(x_{l})\right]\mathcal{L}_{j}\left(x\right)\mathcal{L}_{l}\left(x\right).$
Since for $k,l=1,\ldots,m$, we have
$\mathbb{E}\left[H_{1}(x_{j})H_{1}(x_{l})\right]$ $=\mathbb{E}\left[\left(\mathds{1}_{\left\{\hat{\phi}_{1}\leq x_{j}\right\}}-\mathbb{P}\left(\hat{\phi}_{1}\leq x_{j}\right)\right)\left(\mathds{1}_{\left\{\hat{\phi}_{1}\leq x_{l}\right\}}-\mathbb{P}\left(\hat{\phi}_{1}\leq x_{l}\right)\right)\right]$
$=\mathbb{P}\left(\hat{\phi}_{1}\leq\min\left(x_{j},x_{l}\right)\right)-\mathbb{P}\left(\hat{\phi}_{1}\leq x_{j}\right)\mathbb{P}\left(\hat{\phi}_{1}\leq x_{l}\right)$
$=\min\left(\mathbb{P}\left(\hat{\phi}_{1}\leq x_{j}\right),\mathbb{P}\left(\hat{\phi}_{1}\leq x_{l}\right)\right)-\mathbb{P}\left(\hat{\phi}_{1}\leq x_{j}\right)\mathbb{P}\left(\hat{\phi}_{1}\leq x_{l}\right),$
and using again the strong consistency of $\hat{\phi}_{i}$ for each $i=1,\ldots,N$, we obtain as $n,N\to\infty$,
$\lim\limits_{n,N\to\infty}\mathbb{E}\left[H_{1}(x_{j})H_{1}(x_{l})\right]=\min\left(F(x_{j}),F(x_{l})\right)-F(x_{j})F(x_{l})\mathcal{L}_{j}\left(x\right)\mathcal{L}_{l}\left(x\right),$
which implies that
$\lim\limits_{n,N\to\infty}\mathbb{E}\left(Z_{1,N}^{2}\right)$ $=$ $\sum\limits_{k,l=1}^{m}\left[\min\left(F(x_{j}),F(x_{l})\right)-F(x_{j})F(x_{l})\right]\mathcal{L}_{j}\left(x\right)\mathcal{L}_{l}\left(x\right)$ (7.16)
@@
$\leq$ $\left|\sum\limits_{j=1}^{m}\left(x_{j}-x\right)^{2}\mathcal{L}_{j}\left(x\right)\right|^{1/2}\left|\left(\sum\limits_{j=1}^{m}\mathcal{L}_{j}\left(x\right)^{3}\right)\right|^{1/2}$
$\leq$ $\left[\left(\frac{\pi}{m^{2}}+o\left(m^{-2}\right)\right)S_{m}(x)\right]^{1/2}.$
@@
$S_{m}(x)\leq\dfrac{4}{\pi^{2}}\ln^{2}\left(m+1\right)+1+\dfrac{4}{\pi}\ln\left(m+1\right),$
We then conclude that $J_{m}(x)=O(m^{-1/2})$. Substituting this term in (7.17), we obtain
$\sum\limits_{j=1}^{m}F(x_{j})\mathcal{L}_{j}^{2}\left(x\right)=F(x)S_{m}(x)+O\left(m^{-1/2}\right).$

### v4

@@ 7.7 Proof of Proposition 5
For each $i=1,\ldots,N$, the random effect is expressed by $\phi_{i}=\theta_{i}+\frac{1}{2}\sigma^{2}.$ Then, by a simple plug-in, an estimator of $\phi_{i}$ is given by $\widehat{\phi}_{i}:=\widehat{\theta}_{i}+\frac{1}{2}\widehat{\sigma}^{2},$ where $\theta_{i}$ is given by (3.2) and $\widehat{\sigma}^{2}$ is as expressed in (3.11). From Proposition 1, we already know that $\widehat{\theta}_{i}\xrightarrow{a.s}\theta_{i}$, as $n\to\infty$ and from Proposition 4, $\widehat{\sigma}^{2}\xrightarrow{a.s.}\sigma^{2},$ as $n,N\to\infty$ sequentially. Combining the two results, we conclude that as $n,N\to\infty$ sequentially
$\widehat{\theta}_{i}+\frac{1}{2}\widehat{\sigma}^{2}\xrightarrow{a.s.}\theta_{i}+\frac{1}{2}\sigma^{2}.$
@@
Now we proceed to prove the asymptotic normality of $\widehat{\phi}_{i}$ stated in the second assertion of Proposition 5. We first decompose the estimation error for each $i$, as follows
@@
From Proposition 1, we have $\widehat{\theta}_{i}-\theta_{i}=O_{\mathbb{P}}(n^{H-1})$. Because we let $n\to\infty$ first, for any fixed $N$ we can choose $n$ $n=n_{N}\to\infty$ that grows sufficiently fast (e.g. $n_{N}=N^{2}$), we have
[…]
since $H<\tfrac{3}{4}$ implies $n_{N}^{H-1}$ decays faster than $N^{-1/2}$ for such a choice of $n_{N}$. Thus, under the sequential regime,
$\sqrt{N}\bigl(\widehat{\theta}_{i}-\theta_{i}\bigr)\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}0.$ (7.13)
Moreover, under the same sequential regime, Theorem 1 implies that
$\sqrt{N}\bigl(\widehat{\sigma}^{2}-\sigma^{2}\bigr)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\!\left(0,\;v_{\sigma}^{2}\right),$ (7.14)
where $v_{\sigma}^{2}$ is the asymptotic variance of $\widehat{\sigma}^{2}$.Multiplying (7.12) by $\sqrt{N}$ and combining (7.13) with (7.14), Slutsky’s theorem yields
$\sqrt{N}\bigl(\widehat{\phi}_{i}-\phi_{i}\bigr)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\!\left(0,\;\tfrac{1}{4}v_{\sigma}^{2}\right).$
This completes the proof of the second assertion.
The remainder of this section is devoted to the proofs of the asymptotic properties of the distribution function estimator $\widehat{F}_{m,n,N}$. To that aim, we introduce the following lemma which will be needed later.
@@
The result presented in the next lemma allows us to study the asymptotic properties of $\widehat{F}_{m,n,N}$.
Lemma 3.
Assume that (A1)-(A3) hold and that $n,N\to\infty$ sequentially. Define
$Z_{i,N}:=\sum_{j=1}^{m}\Big(\mathds{1}_{\{\widehat{\phi}_{i}\leq x_{j}\}}-\mathbb{P}(\widehat{\phi}_{1}\leq x_{j})\Big)\mathcal{L}_{j}(x),\quad i=1,\ldots,N.$ (7.15)
Then
$\sup_{1\leq i\neq k\leq N}\left|\mathrm{Cov}\!\left(Z_{i,N},Z_{k,N}\right)\right|=O\!\left(N^{-1}\right),\qquad\text{as }n,N\to\infty\,\text{ sequentially.}$
Proof of Lemma 3
For fixed $i\neq k$, we have
$Z_{i,N}=\sum_{j=1}^{m}\Big(\mathds{1}_{\{\hat{\phi}_{i}\leq x_{j}\}}-\mathbb{P}(\hat{\phi}_{1}\leq x_{j})\Big)L_{j}(x),\qquad Z_{k,N}=\sum_{j=1}^{m}\Big(\mathds{1}_{\{\hat{\phi}_{k}\leq x_{j}\}}-\mathbb{P}(\hat{\phi}_{1}\leq x_{j})\Big)L_{j}(x).$
Hence
$\mathrm{Cov}(Z_{i,N},Z_{k,N})=\sum_{j=1}^{m}\sum_{\ell=1}^{m}L_{j}(x)L_{\ell}(x)\,\mathrm{Cov}\!\left(I_{i,j},I_{k,\ell}\right),$ (7.16)
where $I_{i,j}:=\mathds{1}_{\{\hat{\phi}_{i}\leq x_{j}\}}$.
Recall that $\hat{\phi}_{i}=\hat{\theta}_{i}+\tfrac{1}{2}\hat{\sigma}^{2}$ and denote $\delta:=\tfrac{1}{2}(\hat{\sigma}^{2}-\sigma^{2})$. Since the estimators $\hat{\theta}_{i}$ and $\hat{\theta}_{k}$ are independent, the only source of dependence between $\hat{\phi}_{i}$ and $\hat{\phi}_{k}$ is the common term $\delta$. Hence $I_{i,j}$ and $I_{k,\ell}$ are conditionally independent given $\delta$, i.e. $\mathrm{Cov}(I_{i,j},I_{k,\ell}\mid\delta)=0$. By the law of total covariance,
$\mathrm{Cov}(I_{i,j},I_{k,\ell})=\mathbb{E}\!\left[\mathrm{Cov}(I_{i,j},I_{k,\ell}\mid\delta)\right]+\mathrm{Cov}\!\left(\mathbb{E}[I_{i,j}\mid\delta],\,\mathbb{E}[I_{k,\ell}\mid\delta]\right)=\mathrm{Cov}\!\big(g_{j}(\delta),\,g_{\ell}(\delta)\big),$ (7.17)
where $g_{j}(\delta):=\mathbb{E}[I_{1,j}\mid\delta].$ Conditionally on $\delta$, $\hat{\phi}_{1}=\phi_{1}+\varepsilon_{1}+\delta$ with $\varepsilon_{1}=\hat{\theta}_{1}-\theta_{1}$. From Proposition 1, $\varepsilon_{1}=O_{\mathbb{P}}(n^{H-1})$ and, under the sequential regime $n\to\infty$ first, $\mathbb{E}|\varepsilon_{1}|\to 0$. Consequently,
$g_{j}(\delta)=\mathbb{P}(\phi_{1}+\varepsilon_{1}\leq x_{j}-\delta\mid\delta)=F(x_{j}-\delta)+\eta_{j},$
where $|\eta_{j}|\leq L\,\mathbb{E}|\varepsilon_{1}|$ for some Lipschitz constant $L$ since $F$ is $C^{1}$ by (A2); thus $\eta_{j}=o(1)$ uniformly in $j$. For some $\xi_{j}$ between $x_{j}-\delta$ and $x_{j}$, a Taylor-Young expansion of $F$ gives
$F(x_{j}-\delta)=F(x_{j})-f(x_{j})\delta+\tfrac{1}{2}f^{\prime}(\xi_{j})\delta^{2}.$
Hence
$g_{j}(\delta)=F(x_{j})-f(x_{j})\delta+O(\delta^{2})+o(1).$ (7.18)
Combining (7.17) with (7.18) and noting that $F(x_{j})$ is deterministic, we obtain
$\mathrm{Cov}(I_{i,j},I_{k,\ell})=f(x_{j})f(x_{\ell})\mathrm{Var}(\delta)+O\!\big(\mathbb{E}|\delta|^{3}\big)+o(1).$
Theorem 1 gives $\sqrt{N}\,\delta\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,\tfrac{1}{4}v_{\sigma}^{2})$, hence $\mathrm{Var}(\delta)=O(N^{-1})$ and $\mathbb{E}|\delta|^{3}=O(N^{-3/2})$. Moreover, $f$ is bounded by (A3). Therefore, uniformly in $i\neq k$ and $j,\ell$,
$\big|\mathrm{Cov}(I_{i,j},I_{k,\ell})\big|\leq C\,N^{-1}+o(1).$ (7.19)
Inserting (7.19) into (7.16) yields
$\big|\mathrm{Cov}(Z_{i,N},Z_{k,N})\big|\leq\big(C\,N^{-1}+o(1)\big)\sum_{j=1}^{m}\sum_{\ell=1}^{m}|L_{j}(x)L_{\ell}(x)|.$
For Chebyshev nodes, the Lebesgue constant satisfies $\sup_{x\in[-1,1]}\sum_{j=1}^{m}|L_{j}(x)|=O(\log m)$; consequently
$\sum_{j=1}^{m}\sum_{\ell=1}^{m}|L_{j}(x)L_{\ell}(x)|\leq\Big(\sum_{j=1}^{m}|L_{j}(x)|\Big)^{2}=O((\log m)^{2}).$
Hence
$\big|\mathrm{Cov}(Z_{i,N},Z_{k,N})\big|\leq C\,\frac{(\log m_{N})^{2}}{N}+o(1).$ (7.20)
Because $m=m_{N}$ is chosen to satisfy the growth conditions of Theorems 24 and Proposition 7, we have $(\log m_{N})^{2}=o(N)$ in all relevant asymptotic regimes. Therefore the right-hand side of (7.20) is $O(N^{-1})$, and we finally obtain
$\sup_{1\leq i\neq k\leq N}\big|\mathrm{Cov}(Z_{i,N},Z_{k,N})\big|=O(N^{-1}),$
which completes the proof of Lemma 3.
@@ 7.8 Proof of Proposition 6
For all $x\in[-1,1]$, the expectation of the estimator $\widehat{F}_{m,n,N}$ is given by
$\mathbb{E}\left(\widehat{F}_{m,n,N}(x)\right)$ $=\sum\limits_{j=1}^{m}\frac{1}{N}\sum\limits_{i=1}^{N}\mathbb{E}\left(\mathds{1}_{\left\{\widehat{\phi}_{i}\leq x_{j}\right\}}\right)\mathcal{L}_{j}(x)$
$=\sum\limits_{j=1}^{m}\mathbb{P}\left(\widehat{\phi}_{1}\leq x_{j}\right)\mathcal{L}_{j}(x).$
@@
$\underset{n,N\to\infty}{\lim}\mathbb{P}\left(\widehat{\phi}_{i}\leq x_{j}\right)=\mathbb{P}\left(\phi_{i}\leq x_{j}\right).$ (7.21)
@@
$\lim\limits_{n,N\to\infty}\mathbb{E}\left(\widehat{F}_{m,n,N}(x)\right)=\sum\limits_{j=1}^{m}\mathbb{P}\left(\phi_{1}\leq x_{j}\right)\mathcal{L}_{j}(x)=\sum\limits_{j=1}^{m}F(x_{j})\mathcal{L}_{j}(x).$ (7.22)
@@
$\lim\limits_{n,N\to\infty}\mathbb{E}\left(\widehat{F}_{m,n,N}(x)\right)$ $=F(x)+f(x)\sum\limits_{j=1}^{m}\left(x_{j}-x\right)\mathcal{L}_{j}(x)+\frac{1}{2}f^{\prime}(x)\sum\limits_{j=1}^{m}\left(x_{j}-x\right)^{2}\mathcal{L}_{j}(x)$
@@
$\lim\limits_{n,N\to\infty}\mathbb{E}\left(\widehat{F}_{m,n,N}(x)\right)=F(x)+\frac{\pi}{m^{2}}T_{m}(x)\left[\frac{1}{4}\left(x-1\right)f^{\prime}(x)-\frac{1}{2}f(x)\right]+O(m^{-2}),$
@@
Let us now focus on calculating the variance of our estimator presented by the second assertion of Proposition 6. For all $x\in[-1,1]$,
$\widehat{F}_{m,n,N}(x)-\mathbb{E}\left(\widehat{F}_{m,n,N}(x)\right)$ $=\sum_{j=1}^{m}\!\left(\widehat{F}_{N}(x_{j})-\mathbb{P}(\widehat{\phi}_{1}\leq x_{j})\right)\mathcal{L}_{j}(x)$
$=\frac{1}{N}\sum_{i=1}^{N}Z_{i,N},$
where $Z_{i,N}$ is giving by (7.15) Clearly $\mathbb{E}[Z_{i,N}]=0$ and the variables $Z_{i,N}$ are identically distributed. However, since the estimators $\widehat{\phi}_{i}$ share the common plug–in estimator $\widehat{\sigma}^{2}$, the variables $Z_{i,N}$ are not independent. We therefore use the variance decomposition
$\mathrm{Var}\!\left(\widehat{F}_{m,n,N}(x)\right)$ $=$ $\mathrm{Var}\!\left(\frac{1}{N}\sum_{i=1}^{N}Z_{i,N}\right)$ (7.24)
$=$ $\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}(Z_{i,N})+\frac{2}{N^{2}}\sum_{1\leq i<k\leq N}\mathrm{Cov}(Z_{i,N},Z_{k,N}).$
Since the $Z_{i,N}$ are identically distributed,
$\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}(Z_{i,N})=\frac{1}{N}\mathbb{E}(Z_{1,N}^{2}).$
Define $H_{1}(x_{j}):=\mathds{1}_{\{\widehat{\phi}_{1}\leq x_{j}\}}-\mathbb{P}(\widehat{\phi}_{1}\leq x_{j})$. Then
$\mathbb{E}(Z_{1,N}^{2})$ $=\sum_{j,l=1}^{m}\mathbb{E}\!\left[H_{1}(x_{j})H_{1}(x_{l})\right]\mathcal{L}_{j}(x)\mathcal{L}_{l}(x),$
with
$\mathbb{E}\!\left[H_{1}(x_{j})H_{1}(x_{l})\right]$ $=$ $\mathbb{P}\left(\widehat{\phi}_{1}\leq\min\left(x_{j},x_{l}\right)\right)-\mathbb{P}\left(\widehat{\phi}_{1}\leq x_{j}\right)\mathbb{P}\left(\widehat{\phi}_{1}\leq x_{l}\right)$
$=$ $\min\left(\mathbb{P}\left(\widehat{\phi}_{1}\leq x_{j}\right),\mathbb{P}\left(\widehat{\phi}_{1}\leq x_{l}\right)\right)-\mathbb{P}\left(\widehat{\phi}_{1}\leq x_{j}\right)\mathbb{P}\left(\widehat{\phi}_{1}\leq x_{l}\right)$
Using the Portmanteau theorem again, we obtain
$\lim_{\begin{subarray}{c}n,N\to\infty\end{subarray}}\mathbb{E}\!\left[H_{1}(x_{j})H_{1}(x_{l})\right]=\min(F(x_{j}),F(x_{l}))-F(x_{j})F(x_{l}).$
Hence,
$\lim_{\begin{subarray}{c}n,N\to\infty\end{subarray}}\mathbb{E}(Z_{1,N}^{2})$ $=$ $\sum_{j,l=1}^{m}\left[\min(F(x_{j}),F(x_{l}))-F(x_{j})F(x_{l})\right]\mathcal{L}_{j}(x)\mathcal{L}_{l}(x)$ (7.25)
@@
$\leq$ $\left|\sum\limits_{j=1}^{m}\left(x_{j}-x\right)^{2}\mathcal{L}_{j}\left(x\right)\right|^{\frac{1}{2}}\left|\left(\sum\limits_{j=1}^{m}\mathcal{L}_{j}\left(x\right)^{3}\right)\right|^{\frac{1}{2}}$
$\leq$ $\left[\left(\frac{\pi}{m^{2}}+o\left(m^{-2}\right)\right)S_{m}(x)\right]^{\frac{1}{2}}.$
@@
$S_{m}(x)\leq\dfrac{4}{\pi^{2}}\ln^{2}\left(m+1\right)+1+\dfrac{4}{\pi}\ln\left(m+1\right).$
We then conclude that $J_{m}(x)=O(m^{-\frac{1}{2}})$. Substituting this term in (7.26), we obtain
$\sum\limits_{j=1}^{m}F(x_{j})\mathcal{L}_{j}^{2}\left(x\right)=F(x)S_{m}(x)+O\left(m^{-\frac{1}{2}}\right).$